Two trucks traveled along the same road one at 420 miles per hour and the other at 516 after how many hours did they reach 4680 miles apart?
Answer by checkley77(7172)

Assuming they started from the same place @ the same time then:
516t-420t=4,680
96t=4,680
t=4,680/96
t=48.75 hours after they leave they will be 4,680 mileas apart.
Proof:
516*48.75-420*48.75=4,680
25,155-20,475=4,680
4,680=4,680